# Modeling when and how physics PhD students search for a research group: the role of interests and prior research experiences in timely group integration

### v2

@@ Abstract
Studying the factors that influence the quality of physics PhD students’ doctoral experiences, especially those that motivate them to stay or leave their programs, is critical for providing them with more holistic and equitable support. Prior literature on doctoral attrition has found that students with clear research interests who establish an advisor-advisee relationship early in their graduate careers are most likely to persist [1]. However, these trends have not been investigated in the context of physics, and the underlying reasons for why these characteristics are associated with leaving remain unstudied. Using semi-structured interviews with 40 first and second year physics PhD students, we construct a model describing the characteristic pathways that physics PhD students take while evaluating interest congruence of prospective research groups. We show how access to undergraduate research and other formative experiences helped some students narrow their interests and look for research groups before arriving to graduate school. In turn, these students reported fewer difficulties finding a group than students whose search for an advisor took place during the first year of their PhD. Lastly, we identify two characteristic types of students at a higher risk of leaving their programs: students who enter graduate school with broad interests and struggle to find a group, and students who join a research group early based on research interest alone and subsequently encounter issues with a negative mentoring relationship. This work serves as a major step toward creating a comprehensive model of how PhD students find a research group, and opens the door for future work to investigate how factors such as group culture and working environment impact the search process.
@@ I Introduction and Background
Graduate school is known to be a complex and intense environment for first-year PhD students. Adapting to the norms and expectations of a new and unfamiliar department can be difficult, as can the process of transitioning into the role of professional researcher. During this critical period, finding the right research advisor is widely acknowledged to be one of the most impactful events for students on the path to a PhD [2, 3]. Finding an effective research mentor and group provides the basis for students to build specialized research skills [4, 5], embrace their identities as scientists [6, 7], and developing a network of support among their labmates [8, 9]. On the other hand, poor advising relationships negatively affect PhD students’ intellectual development, mental health, and willingness to pursue an academic career after graduation [10, 11, 12]. For some students, a negative advising relationship is one of the main factors that motivates them to leave their programs [1, 13, 14, 3, 15, 16].
@@
Studying the factors that influence the quality of PhD students’ experience and motivate them to stay or leave their programs is an important goal for researchers across graduate education [17, 18, 19, 20, 21]. Better understanding why doctoral students leave is particularly critical for creating a more diverse and equitable graduate landscape, as attrition from PhD programs disproportionately affects traditionally underrepresented students [22, 23, 24]. In the context of physics, the retention rate of physics PhD students is estimated to be approximately 50% [25, 22]. The issue of high attrition is important for both students and the institutions they attend. For students, leaving a PhD program may mean grappling with feelings of failure and disappointment, which can have adverse effects on their emotional and physical well-being [1]. Graduate departments must invest future resources into recruiting and supporting new students [26].
@@
In a seminal multidisciplinary study of doctoral attrition, Lovitts’ study found that a combination of students’ research interests, proactivity in seeking out a research group, and closeness of interaction with prospective advisors prior to graduate school were all predictive of students’ likelihood of PhD completion. She notes that “few noncompleters enter graduate school knowing who they want to work with,” and that “having interests that are close to a faculty member’s and establishing the adviser-advisee relationship early in the graduate career based on firsthand knowledge and interaction” contribute most to successful PhD completion. Thus, Lovitts contends that students with strong research interests who search for and join a group early are most likely to persist. However, the underlying reasons driving these trends remain unclear. For example, it is not obvious that having broad research interests should correlate with a lower likelihood of PhD completion; on the contrary, intellectual curiosity and a willingness to explore new topics are traits often valued in science. In fact, physics was one discipline not included in Lovitts’ study, leaving it unknown whether these trends apply to physics students as well. To fill this gap, this study seeks to characterize physics PhD students’ research interests as they enter graduate school, and to better understand the influence that these interests have on their search for a research group. In doing so, we also seek to better understand how these predictors of attrition (strength of interest, time students look for and join research groups) are linked, as well as uncover the underlying factors driving these trends.
@@
Little is known about the extent to which physics PhD students know what research they want to pursue upon entering graduate school. However, studies by Maher et al. in the context of biology PhD programs have shown that biology graduate students place a high emphasis on finding a group with congruent research interests, despite often being unsure of what research they wanted to pursue. Biology PhD students commonly hoped to explore their interests while in graduate school [28, 27], and interest exploration in biology is encouraged through highly formalized lab rotation systems. The authors found these rotations to benefit students trying to narrow their interests. Such formal requirements are comparatively rare in physics graduate programs though [29, 30]. Rather, as detailed in our previous work, physics PhD students often sought more structure for meeting groups, which may harm students with broader interests if they encounter more difficulties looking for a group [31].
@@
Our prior work also demonstrated that physics graduate students characterize finding the right research group as a highly important and impactful part of their graduate careers, but often encountered a variety of difficulties in their group search [31] . These included difficulties understanding the research that prospective groups were doing, struggling to connect with faculty and graduate students, and lacking time to search due to coursework. Moreover, we illustrated how students who struggled to find a group tended to report a lower sense of belonging in their programs, which led several of them to consider leaving their programs. On the other hand, successfully joining a group often granted students a greater sense of community and purpose in their programs that motivated them to stay. These results are supported by prior research that has shown how becoming a part of a research group can give STEM students opportunities to develop their identities as scientists [6, 32] while increasing their sense of belonging to their research community [33, 9, 34, 35, 36].
@@
Access to a research group, and all of the benefits that are known accompany it, may depend on how efficiently students are able to navigate the search process. Most attrition in physics PhD programs occurs in the first two years of the program [42], but research shows that students begin to think about leaving begin much earlier than that. Lovitts revealed that students who leave after their first year of graduate school typically begin contemplating their decision to leave in the middle of their second semester. This timeframe coincides with when many students are likely in the midst of attempting to find a research group, which may therefore play a critical role in their decision to stay or leave. However, tracking when students join and leave research groups is not a common metric for departments to record, which leaves it unclear when students actually take this critical step in their graduate careers. Hidden inequities are likely to go unnoticed in graduate education when student outcomes and progress are poorly tracked [1]. Hence, better understanding the factors that influence when students join a research group may yield insight into systemic inequities inherent in the group search process.
For instance, ample research has illustrated that physics students’ interests are most informed by their backgrounds and learning experiences [43, 44, 45], and that access to undergraduate research experiences are particularly critical [34, 36, 46]. Looking for groups takes time and effort, and students often described having trouble having enough time to evaluate groups while inundated with coursework and teaching responsibilities. Thus, students whose backgrounds offer them more opportunities to narrow their research interests and make connections with graduate faculty before entering graduate school may be better positioned to more quickly begin research. Moreover, students’ backgrounds and undergraduate experiences strongly impact their navigation of the group search process. In our previous work, we identified a number of advantages that one student enjoyed while navigating the search process: attending an elite undergraduate institution provided a leg up in admissions, an REU at his top-choice institution gave insight into which labs he might want to join, and advice from his undergraduate advisor helped him evaluate prospective advisors [47]. Meanwhile, another student described difficulties navigating graduate school, and expected significantly more academic advising from her graduate school to help her through her first year. This student struggled to find a research lab, which negatively influenced her sense of belonging in the program.
Our study is guided by the following research questions:
@@
Sections IV.1, IV.2, and IV.4 attend to answering RQ1. In Section IV.1, we outline when students in our sample found their research groups and identify a subset of students in our sample who committed to a research group particularly early in their PhD timelines. We also illustrate a number of benefits that these students associated with doing so. Although we acknowledge that group culture plays an important role in the group search process (see Figure 1), RQ2 highlights that this study is primarily focused on how students find groups with congruent research interests RQ2 is discussed in Section IV.2, and demonstrates that physics PhD students enter graduate school with drastically different levels of certainty regarding which research they wanted to pursue, and then discussed how this correlates with when students join research groups. Then, in Section IV.4, we address RQ3 by investigating when students started looking into research groups, and shows how students who engage in the search process earlier tend to join a group more quickly. Throughout both of these sections, we provide evidence to highlight how students’ access to resources and experiences impacted their search process. Lastly, in Section IV.5, we show how the frequency and types of concerns and difficulties that students experience during their search process vary depending on when they joined a research group. This analysis exposes some of the costs and benefits associated with joining a research group early, and culminates in a model that identifies two characteristic types of students at a higher risk of leaving their programs.
@@ II Conceptual Framing
We leveraged insights from two theoretical frameworks while designing the current study, following the example of Maher et al. [27]. This section briefly describes these frameworks and how they informed our methodological approach. First, we discuss person-environment fit, which has recently been suggested as a useful theoretical lens through which to examine doctoral education [48]. We then discuss socialization theory, which is the most prevalent framework used in studies of US doctoral education [49].
@@
The final stage of socialization is internalizing the role of an independent scholar and integrating their new identity into their existing ones. Socialization frameworks are often used as a lens to examine attrition from doctoral programs, and poor socialization has been linked to students’ decisions to leave graduate school [1, 67]. These findings highlight how inadequate structural support for students, rather than perceived shortcomings in individual students’ behaviors and characteristics, can lead to their early departure. However, socialization has sometimes been criticized for downplaying the role of student agency as they move through their programs [68, 66], leading to more recent iterations of the framework to more explicitly include students’ role in advancing their own socialization [69, 62].
@@
Table 1: Gender and race data for the $N=40$ interviewees in the sample Demographic information was gathered using a Qualtrics survey prior to each interview. Surveys were fixed-choice, but contained a space for respondents to input answers outside the choices provided.
@@ III.1 Methodological approach and interview protocol
We sought out methodological approaches that would allow us to explore the role of student agency and structural constraints in the group search process. We therefore drew inspiration from cognitive task analysis (CTA) methods [70, 71] and Dervin’s sense-making method [72, 73] in out effort to characterize the process by which PhD physics students search for a research group. CTA methods seek to elucidate the cognitive processes underlying individual behavior, particularly regarding how individuals evaluate and respond to their external environment [74, 75, 76]. Similarly, sense-making operates under the assumption that humans move in and out of structures that influence their behavior, sometimes making sense of them, resisting them, or creating new structures altogether [77]. Both methodologies are designed to elicit detailed descriptions of interviewees’ thoughts and actions as they recount how they progressed toward a goal, focusing on understanding specific life experiences in great depth.
@@
Our semi-structured interview protocol asked students to construct a timeline of steps they took while searching for a research group (Stage 1). Students were free to start their timeline at any point, but most commonly they began during their junior or senior years of undergraduate study. Timelines often consisted of four or five main steps. For each step, we asked students about any major questions and concerns they had at each step of their timeline (Stage 2). This included anything they wanted to find out, were confused about, worried about, or were just curious about. Many responses to this part of the protocol revolved around students’ internal deliberations regarding what research they wanted to pursue in graduate school, as well as their efforts to gather information about prospective groups’ research. Students also often commented on general difficulties they experienced during their search, as well as concerns about their future in graduate school. Lastly, we asked students to identify any sources of help that allowed them to resolve their question or concern, as well as any obstacles that hurt their ability to move forward (Stage 3). This third stage allowed us to understand the thoughts, actions, and events that helped or hurt students’ ability to navigate their search for a group. The analysis presented here primarily focuses on several of the major questions and concerns that students discussed in Stage 2, but some supporting evidence was drawn from Stage 3 of the protocol focusing on sources of helps and hurts. We leveraged the data from Stage 1 to identify when each event occurred in their timeline. The full interview protocol used in this study is available in the Supplemental Material [78].
@@
When asked to describe the major steps in their timeline, we also asked students to identify when in their academic careers those events occurred. Students typically described when events occurred in terms of academic periods. Using this information, we were able to identify each excerpt as occurring at one of seven points in time: Before Senior year, Senior year, Senior summer, 1st semester, 2nd semester, 1st year summer, and 2nd year. Senior year refers to students’ final year of undergraduate study, whereas 1st and 2nd semester refer to the first and second semesters in their graduate program. We use this language in the results section IV to indicate when in the student’s timeline they said it.
@@ III.2 Recruitment and participant sample
Since this project is conducted with the support of the Inclusive Graduate Education Network (IGEN), a partner of the American Physical Society (APS), one major goal of this work is to improve diversity and inclusion across physics graduate education. One of the leading efforts toward that end is the APS Bridge Program, which is a post-baccalaureate program designed to increasing the number of PhDs earned by underrepresented students in physics [80]. Hence, we also sent our recruitment information directly to APS Bridge students currently in their PhD programs to help ensure our data represented their experiences as well.
The sample of students in this analysis consists of 40 students representing 13 institutions; 5 students were from the Bridge program, all from different institutions. A fixed-choice demographic survey was administered prior to each interview via Qualtrics, but each response field contained a space for respondents to input answers outside the choices provided. Based on the results of the survey, $N=22$ interviewees identified as women, $N=15$ identified as men, and $N=3$ identified as non-binary or gender fluid. $N=21$ were in their first year of physics graduate school while $N=19$ were in their second year. $N=18$ identified as White/Caucasian, $N=9$ as Asian, $N=8$ as Hispanic/Latinx or Spanish origin, $N=2$ as Black or African American, $N=1$ as North African, and $N=2$ from multiple races. $N=32$ students were from the US while $N=8$ were non-US students. Table II offers a more detailed demographic breakdown. Interviews were conducted over Zoom and took place from Fall 2022 to Spring 2023. All names used throughout the paper are pseudonyms.

### v3

@@ Abstract
Studying the factors that influence the quality of physics PhD students’ doctoral experiences, especially those that motivate them to stay or leave their programs, is critical for providing them with more holistic and equitable support. Prior literature on doctoral attrition has found that students with clear research interests who establish an advisor-advisee relationship early in their graduate careers are most likely to persist. However, these trends have not been investigated in the context of physics, and the underlying reasons for why these characteristics are associated with leaving remain unstudied. Using semi-structured interviews with 40 first and second year physics PhD students, we construct a model describing the characteristic pathways that physics PhD students take while evaluating interest congruence of prospective research groups. We show how access to undergraduate research and other formative experiences helped some students narrow their interests and look for research groups before arriving to graduate school. In turn, these students reported fewer difficulties finding a group than students whose search for an advisor took place during the first year of their PhD. Lastly, we identify two characteristic types of students at a higher risk of leaving their programs: students who enter graduate school with broad interests and struggle to find a group, and students who join a research group early based on research interest alone and subsequently encounter issues with a negative mentoring relationship. This work serves as a major step toward creating a comprehensive model of how PhD students find a research group, and opens the door for future work to investigate how factors such as group culture and working environment impact the search process.
@@ I Introduction and Background
Graduate school is known to be a complex and intense environment for first-year PhD students. Adapting to the norms and expectations of a new and unfamiliar department can be difficult, as is the process of transitioning into the role of professional researcher. During this critical period, finding the right research advisor is widely acknowledged to be one of the most impactful events for students on the path to a PhD [1, 2]. Finding an effective research mentor and group provides the basis for students to build specialized research skills [3, 4], embrace their identities as scientists [5, 6], and develop a network of support among their labmates [7, 8]. On the other hand, poor advising relationships negatively affect PhD students’ intellectual development, mental health, and willingness to pursue an academic career after graduation [9, 10, 11]. For some students, a negative advising relationship is one of the main factors that motivates them to leave their programs [12, 13, 14, 2, 15, 16].
@@
Studying the factors that influence the quality of PhD students’ experiences and motivate them to stay or leave their programs is an important goal for researchers across graduate education [17, 18, 19, 20, 21]. Better understanding why doctoral students leave is particularly critical for creating a more diverse and equitable graduate landscape, as attrition from PhD programs disproportionately affects traditionally underrepresented students [22, 23, 24]. In the context of physics, the retention rate of PhD students is estimated to be approximately 50% [25, 22]. For students, leaving a PhD program may mean grappling with feelings of failure and disappointment, which can have adverse effects on their emotional and physical well-being [12]. Graduate departments are not left unscathed by high rates of attrition either, as they must invest future resources into recruiting and supporting new students [26].
Most attrition from physics PhD programs occurs in the first two years [27], but research shows that students begin to think about leaving long before they actually do so. Students who leave after their first year of graduate school typically begin contemplating their decision to leave by their second semester [12]. This timeframe coincides with when many students are likely in the midst of attempting to find a research group, and their success in that process may impact their decision to stay or leave. However, tracking when students join and leave research groups is not a common metric for departments to record, which leaves it unclear when students actually take this critical step in their graduate careers. Hidden inequities are most likely to go unnoticed in graduate education when student outcomes and progress are poorly tracked [12]; thus, better understanding the factors that influence when students join a research group may yield insight into systemic inequities inherent in the group search process.
@@
In a seminal multidisciplinary study of doctoral attrition, Lovitts found that a combination of students’ research interests, proactivity in seeking out a research group, and closeness of interaction with prospective advisors prior to graduate school were all predictive of students’ likelihood of PhD completion. She noted that “few noncompleters enter graduate school knowing who they want to work with,” and that “having interests that are close to a faculty member’s and establishing the advisor-advisee relationship early in the graduate career based on firsthand knowledge and interaction” contribute most to successful PhD completion. Thus, Lovitts contends that students with strong research interests who search for and join a group early are most likely to persist.
However, the underlying reasons driving these trends remain unclear. For example, it is not obvious that having broad research interests should correlate with a lower likelihood of PhD completion; on the contrary, intellectual curiosity and a willingness to explore new topics are traits often valued in science. Furthermore, physics was one discipline omitted in Lovitts’ study, leaving it unknown whether these trends apply to physics students as well. To fill this gap, this study seeks to characterize physics PhD students’ research interests as they enter graduate school, and to better understand the influence that these interests have on their search for a research group. In doing so, we also seek to better understand how these previously identified predictors of attrition (strength of interest, time students look for and join research groups) are linked, as well as uncover the underlying factors driving these trends.
@@
Little is known about the extent to which physics PhD students know what research they want to pursue upon entering graduate school. Studies by Maher et al. in the context of biology PhD programs have shown that biology graduate students place a high emphasis on finding a group with congruent research interests, despite often being unsure of exactly what research topic they were interested in pursuing. Biology PhD students commonly hoped to explore their interests while in graduate school [29, 28], and interest exploration in biology is encouraged through highly formalized lab rotation systems. The authors found these rotations benefited students in trying to narrow their interests. Such formal requirements are comparatively rare in physics graduate programs though [30, 31]. Rather, as detailed in our previous work, physics PhD students often sought more structure for meeting groups, which may harm students with broader interests if they encounter more difficulties finding the right fit [32].
@@
That study showed other difficulties that physics students described as well, including trouble understanding the research that prospective groups were doing, struggling to connect with faculty and graduate students, and lacking time to search due to coursework. Moreover, students who struggled to find a group tended to report a lower sense of belonging in their programs during their first year, which led several of them to consider leaving their programs. On the other hand, successfully joining a group often granted students a greater sense of community and purpose in their programs that motivated them to stay. These results are supported by prior research that has shown how becoming a part of a research group can give STEM students opportunities to develop their identities as scientists [5, 33] while increasing their sense of belonging to their research community [34, 8, 35, 36, 37].
@@
Yet access to a research group, and all of the benefits that are known accompany it, may strongly depend on students’ research interests and how efficiently they are able to navigate the search process. Students whose backgrounds offer them more opportunities to narrow their research interests and make connections with graduate faculty before entering graduate school may be better positioned to more quickly begin research. Research on the impacts of undergraduate research experiences has shown that students tend to constrain their options for future work to topics with which they are familiar [43, 44]. Indeed, ample research has illustrated that physics students’ interests are most informed by their backgrounds and learning experiences [45, 46, 47], and access to undergraduate research experiences are particularly critical [35, 37, 48].
Despite the fact that undergraduate research experiences are often critical for students’ interest development, access to undergraduate research is lower for traditionally underrepresented groups in physics, particularly ethnic minorities and first-generation college students [49]. In another previous work, we revealed how access to undergraduate research can have significant repercussions during the group search [50]. We identified a number of advantages that one student enjoyed while navigating the search process: attending an elite undergraduate institution provided a leg up in admissions, an REU at his top-choice institution gave insight into which labs he might want to join, and advice from his undergraduate advisor helped him evaluate prospective advisors. Meanwhile, another student described a number of difficulties navigating graduate school, many of which she attributed to a lack of departmental support for first-generation students. She expected significantly more academic advising from her graduate school to help her through her first year and struggled to find a research lab, which negatively influenced her sense of belonging in the program. Collectively, prior literature suggests relationships between students’ background experiences, their research interests, and when and how they go about searching for a research group in graduate school.
In this study we seek to better understand how physics PhD students find a group with congruent research interests. While we acknowledge that other characteristics (e.g., the group’s culture, work-life balance) are also important to students, the present work focuses on the role of research interests and was guided by the following questions:
@@
Sections IV.1 to IV.4 attend to answering RQ1 and RQ2. In Section IV.1, we outline when students in our sample found their research groups and identify a subset of students in our sample who committed to a research group early in their PhD timelines. We also illustrate a number of benefits that these students associated with doing so. Section IV.2 demonstrates that physics PhD students enter graduate school with drastically different levels of certainty regarding which research they wanted to pursue, and Section IV.3 discusses how this correlates with when students join research groups. Then, Section IV.4 investigates when students look for research groups, showing how students who engage in the search process earlier tend to join a group more quickly. Throughout each of these sections, we provide evidence to highlight how students’ access to resources and experiences during their undergraduate studies impacted their search process. In Section IV.5, we address RQ3 by showing how the frequency and types of concerns and difficulties that students experience during their search process vary depending on when they joined a research group. Lastly, Section IV.6 focuses on RQ4 by closely examining the cases of several students who considered leaving their PhD programs, culminating in a model that identifies two characteristic types of students at a higher risk of leaving.
@@ II Conceptual Framing
We leveraged insights from two theoretical frameworks while designing the current study, following the example of Maher et al. [28]. This section briefly describes these frameworks and how they informed our methodological approach (see Figure 1). First, we discuss person-environment fit, which has recently been suggested as a useful theoretical lens through which to examine doctoral education [51]. We then discuss socialization theory, which is the most prevalent framework used in studies of US doctoral education [52].
@@
The final stage of socialization is internalizing the role of an independent scholar and integrating this new identity into existing ones. Socialization frameworks have been used as a lens to examine attrition from doctoral programs, and are particularly useful for conceptualizing the dual nature of PhD students as both students and professionals. Golde refers to this as an “unusual double socialization” of graduate students [15], who in addition to integrating into their individual department in their role as students must also begin to operate as professional physicists in their role as researchers [70]. Along both of these axes, faculty clearly play a critical role as both instructors and advisors. Critically, poor socialization has been linked to students’ decisions to leave graduate school [12, 71]. These findings highlight how inadequate structural support for students, rather than perceived shortcomings in individual students’ behaviors and characteristics, can lead to early departure. However, socialization has sometimes been criticized for downplaying the role of student agency as they move through their programs [72, 69], leading to more recent iterations of the framework to more explicitly include the student’s role in advancing their own socialization [73, 65].
@@
Table 1: Gender and race data for the $N=40$ interviewees in the sample Demographic information was gathered using a Qualtrics survey prior to each interview. Surveys were fixed-choice, but contained a space for respondents to input answers outside the choices provided. Fixed-choice responses for race/ethnicity questions were drawn from the US Census American Community Survey [74].
@@ III.1 Methodological approach and interview protocol
We sought out methodological approaches that would allow us to explore the role of student agency as well as structural constraints in the group search process. We drew inspiration from cognitive task analysis (CTA) methods [75, 76] and Dervin’s sense-making method [77, 78]. According to Crandall et al. in their handbook on cognitive task analysis, CTA methods are meant to “understand how people think: how their minds work, what they struggle with, and how they manage to perform complex work adeptly” [75]. For example, analogous to the present work on how students “shop” for a research advisor, such methods have been used to study how consumers gather information about products and choose which ones to purchase [79]. One critical aspect of understanding the cognitive processes underlying individual behavior is how individuals evaluate and respond to their external environment [80, 81, 82]. Similarly, sense-making operates under the assumption that humans move in and out of structures that influence their behavior, sometimes making sense of them, resisting them, or creating new structures altogether [83]. Both methodologies are designed to elicit detailed descriptions of interviewees’ thoughts and actions as they recount how they performed a task, focusing on understanding their specific life experiences in great depth.
@@
Our semi-structured interview protocol asked students to construct a timeline of steps they took while searching for a research group (Stage 1). Students were free to start their timeline at any point, but most commonly they began during their junior or senior years of undergraduate study. Timelines often consisted of four or five main steps. For each step, we asked students about any major questions and concerns they had at each step of their timeline (Stage 2). This included anything they wanted to find out, were confused about, worried about, or were just curious about. Many responses to this part of the protocol revolved around students’ internal deliberations regarding what research they wanted to pursue in graduate school, as well as their efforts to gather information about prospective groups’ research. Students also often commented on general difficulties they experienced during their search, as well as concerns about their future in graduate school. Lastly, we asked students to identify any sources of help that allowed them to resolve their question or concern, as well as any obstacles that hurt their ability to move forward (Stage 3). This third stage allowed us to understand the thoughts, actions, and events that helped or hurt students’ ability to navigate their search for a group. The analysis presented here primarily focuses on several of the major questions and concerns that students discussed in Stage 2, but some supporting evidence was drawn from Stage 3 of the protocol focusing on sources of helps and hurts. The full interview protocol used in this study is available in the Supplemental Material [84].
@@
We also leveraged the data from Stage 1 to identify when each event occurred in their academic timeline. Students typically described when events occurred in terms of academic periods. Using this information, we were able to identify each excerpt as occurring at one of seven points in time: Before Senior year, Senior year, Senior summer, 1st semester, 2nd semester, 1st year summer, and 2nd year. Senior year refers to students’ final year of undergraduate study, whereas 1st and 2nd semester refer to the first and second semesters in their graduate program. We use this language in the results section IV to indicate when in the student’s timeline they said it.
@@ III.2 Recruitment and participant sample
Since this project is conducted with the support of the Inclusive Graduate Education Network (IGEN), a partner of the American Physical Society (APS), one major goal of this work is to improve diversity and inclusion across physics graduate education. One of the leading efforts toward that end is the APS Bridge Program, which is a post-baccalaureate program designed to increase the number of PhDs earned by underrepresented students in physics [86]. Hence, we also sent our recruitment information directly to APS Bridge students currently in their PhD programs to help ensure our data represented their experiences as well.
The sample of students in this analysis consists of 40 students representing 13 institutions; 5 students were from the Bridge program, all from different institutions. A fixed-choice demographic survey was administered prior to each interview via Qualtrics, but each response field contained a space for respondents to input answers outside the choices provided. Based on the results of the survey, $N=22$ interviewees identified as women, $N=15$ identified as men, and $N=3$ identified as non-binary or gender fluid. $N=21$ were in their first year of physics graduate school while $N=19$ were in their second year. Fixed-choice responses for race/ethnicity questions were drawn from the US Census American Community Survey [74]; $N=18$ identified as White/Caucasian, $N=9$ as Asian, $N=8$ as Hispanic/Latinx or Spanish origin, $N=2$ as Black or African American, $N=1$ as North African, and $N=2$ from multiple races. $N=32$ students were from the US while $N=8$ were non-US students. Table II offers a more detailed demographic breakdown. Interviews were conducted over Zoom and took place from Fall 2022 to Spring 2023. All names used throughout the paper are pseudonyms.